An Frisbee of mass 175 g (note: 1000g = 1 kg) flies through the air at a speed of 5.5 m/s at a height of 1.7 metres. What is the Frisbees total mechanical energy?
a) 2.9 J
b) 5.6 J
c) 12 J
d) 46 J
1
Expert's answer
2020-05-26T12:43:37-0400
"E=0.5mv^2+mgh=m(gh+0.5v^2)"

"E=0.175((9.8)(1.7)+0.5(5.5)^2)=5.6\ J"

Answer: b) 5.6 J